# VoiceAgentBench: Are Voice Assistants ready for agentic tasks?

### v1

@@ 1 Introduction
Advancements in Large Language Models (LLMs) (Touvron et al., 2023; Grattafiori et al., 2024; Abdin et al., 2025; Guo et al., 2025; Yang et al., 2025b) have enabled the development of intelligent agents capable of reasoning (Wei et al., 2022), planning (Yao et al., 2023), and executing complex, multi-step tasks through interaction with external tools (Qin et al., 2024; Patil et al., 2024) and databases (Gao et al., 2024). These agentic systems have shown strong performance on tasks such as code generation (Rozière et al., 2024; DeepSeek-AI et al., 2024), document question answering (Zhao et al., 2024), and interactive AI applications, highlighting their potential to automate sophisticated workflows. Most existing research, however, focuses on text-based interactions, overlooking speech as a natural and accessible modality. Extending agents to voice is critical for hands-free, conversational control in real-world applications. Current approaches typically depend on external automatic speech recognition (ASR) models, leaving open the fundamental question of how agents can directly process and respond to spoken input without an intermediate step. SpeechLMs [^1] Although existing models and benchmarks cover multiple formats, including audio, speech, and music, our focus here is specifically on human speech tasks; henceforth we use the term SpeechLMs. capable of instruction following and chat-style interactions (Xu et al., 2025; KimiTeam et al., 2025) pave the way toward such agents. By avoiding the sequential transcription overhead of ASR–LLM pipelines, SpeechLMs enable lower latency (Appendix F) and natural real-time interactions. They further demonstrate the ability to interpret spoken commands, maintain context across turns, and generate task-relevant responses in natural language, providing a strong foundation for voice assistants.
@@ 2 Related Work
LLM Agent Benchmarks. Interest in evaluating agentic LLMs has grown with advances in their reasoning and decision making capabilities. ToolBench (Qin et al., 2024) evaluates models’ ability to invoke external APIs across diverse real-world tasks, while ToolQA (Zhuang et al., 2023) assesses LLMs’ use of external tools for question answering via a scalable, automated dataset curation process. Berkeley Function Calling Leaderboard (BFCL) (Patil et al., 2025) emphasizes precise API generation across domains and robustness to both single and multiple function calls, and NESTful (Basu et al., 2025) focuses on nested sequences of API calls, where outputs of one call feed into the next. API-Bank and ToolTalk (Li et al., 2023; Farn & Shin, 2023) target multi-turn, dialogue-driven tool-use scenarios, testing sequential API planning and interaction. Tau-bench (Yao et al., 2025) simulates dynamic conversations with domain-specific tools and policies to evaluate adherence to task rules. AgentHarmBench (Andriushchenko et al., 2025) and DoomArena (Boisvert et al., 2025) focus on safety and adversarial robustness, testing susceptibility to harmful or unsafe actions. Despite this progress for LLMs, no speech benchmark explicitly evaluates SpeechLMs in such realistic, agentic, and safety-critical settings, underscoring the need for specialized evaluation frameworks.
@@
Speech Datasets and Benchmarks. Large-scale datasets such as LibriSpeech (Panayotov et al., 2015), CommonVoice (Ardila et al., 2020), and MuST-C (Di Gangi et al., 2019) have been foundational for advancing automatic speech recognition (ASR) and speech translation (AST). IndicST (Shah et al., 2025) and Lahaja (Javed et al., 2024) extend these tasks to cover diverse Indic speech data. Evaluation suites like SUPERB (wen Yang et al., 2021) and SLUE (Shon et al., 2023) standardize the assessment of tasks such as intent classification, named entity recognition, and slot filling, with IndicSUPERB (Javed et al., 2022) further supporting Indic languages. However, these benchmarks primarily target simpler tasks like transcription, translation, NER and do not assess reasoning or decision-making over spoken content. To address this gap, recent work has begun exploring reasoning in the audio domain. Audio-CoT (Ma et al., 2025) introduces chain-of-thought prompting for structured multistep inference on speech input, while MMAU (Sakshi et al., 2025) provides a large-scale benchmark of 10k audio clips covering 27 reasoning skills, such as temporal reasoning and causal inference, in speech, music, and environmental sounds. AIR-Bench (Yang et al., 2024) and AudioBench (Wang et al., 2024) extend the scope to open-ended instruction following on various types of audio and speech, whereas VoiceBench (Chen et al., 2024a) emphasizes robustness and generalization by converting text instruction into spoken form with real-world noise and speaker variation. More recently, SpeechR (Yang et al., 2025c) directly targets high-level reasoning on speech, focusing on logical deduction, and commonsense problem solving. We also provide an extended discussion of related work on speech models in Appendix B.
@@ 3.2.1 Tool Sourcing
Single Tool, Single Tool with Retrieval, and Parallel Tool Invocation. For these categories, we leverage BFCL (Patil et al., 2025), which provides well-structured functions for such cases. Single Tool Call directly adapts the simple tool subset from BFCL. The Single Tool with Retrieval category uses the multiple tool subset, requiring the model to select the most relevant tool and fill the arguments. The Parallel Tool category leverages the parallel multiple subset, where multiple independent tool calls are invoked simultaneously. In total, we incorporate 458 functions here.
Sequential Dependent Tools. While benchmarks like NESTful (Basu et al., 2025) focus on sequential tool invocation in specialized domains such as mathematics and coding, they do not capture the practical, everyday tasks expected of real-world voice assistants. These tasks often require chaining interdependent tools to complete workflows such as booking a cab, ordering food, or managing payments. To address this, we designed a set of 21 tools across three realistic agents: i) Cab booking, ii) Food ordering, and iii) Payment services. The toolsets are presented in Appendix H.
Multi-Turn Dialogue Tools. For dialogue-based tool invocation, we adopt tools from the Level-2 subset of API-Bank (Li et al., 2023), which are designed to support Retrieval+Call based on multi-turn user interactions. Using this subset, we incorporate a total of 49 tools for this category, enabling evaluation of conversationally grounded, multi-turn tool-calling capabilities.
Safety Evaluation Tools. For evaluating the safety of agentic behavior, we utilize tools from the AgentHarm (Andriushchenko et al., 2025). It encompasses 11 harm categories, including fraud, cybercrime, and harassment. These tasks are designed to assess whether models can refuse harmful agentic requests. By integrating these tools, we enable comprehensive evaluation of an agent’s robustness against adversarial and unsafe tool usage scenarios.
@@ 3.2.2 Text Query & Dialogue Generation
ii). Sequential Dependent Tool Calling (SeqDep): To capture realistic use cases, we design three service agents: cab booking, food ordering, and payments by creating seed data with available tools and expected outputs. Using GPT-4o-mini (Hurst et al., 2024), we generate queries conditioned on tool schemas and dependency chains, ensuring multiple interdependent calls (e.g., booking a cab after retrieving location coordinates or completing a food order with stored address and payment details). This yields 40 complex queries each in English and Hindi, paired with tool-call responses, capturing dependency-driven tasks beyond simple or parallel invocation.
@@ 3.2.3 Model Instructions and Indic Multilingual Extension
To standardize behavior across models, we design category-specific system instructions that direct the model to produce tool calls strictly in Python syntax, following Patil et al. (2025). This prevents free-form or mixed response which cannot be deterministically parsed. To further anchor the model’s behavior, we provide a one-shot example in the instruction that clarifies the output format without imposing strong task-specific biases. For multilingual settings, we additionally instruct the model to generate tool calls exclusively in English, aligning with real-world usage scenarios. Additionally, for safety evaluations we append a refusal prompt in the instructions (examples in Appendix J.1).
We extend VAB to five Indic languages by translating the English Indian context queries into the target languages. For Malayalam, we employ Llama-3.3 70B, while for the remaining languages we use Gemma3 27B, this is based on human eval results shown in (Anonymous, 2025). We complement this with a lightweight quality control pipeline that flags issues such as script mixing and unknown tokens. The validated queries are then fed into the diversity based TTS pipeline.
@@ 3.2.4 Diversity Based TTS Generation
In synthetic speech generation, the absence of real speakers and natural voices necessitates methods that ensure diversity in the generated data, motivating new selection strategies to build robust and representative benchmarks. Following IndicSynth (Sharma et al., 2025), which employed the VoxLingua107 ECAPA-TDNN model (Desplanques et al., 2020), trained on diverse Indic languages and accents, to evaluate the linguistic authenticity of synthetic speech, we adopt ECAPA-TDNN embeddings for our diversity analysis. Adapting diversity principles from word embedding literature, we transfer these ideas to the audio domain using ECAPA-TDNN embeddings. Specifically, we ablate three strategies for selecting maximally diverse audio samples: i) Determinantal Point Processes (DPP) (Wang et al., 2025), ii) Farthest Point Sampling (FPS) adapted from PointNet++ (Qi et al., 2017), and iii) a Density-based Probabilistic Method (Appendix C.1). Diversity is quantified using the mean distance to the nearest selected point (Yang et al., 2025d), a metric that captures coverage of the embedding space. Our evaluation (Appendix C.2) shows that FPS (Algorithm 1) consistently achieves the highest diversity scores on our dataset, establishing it as the most effective strategy. We conduct this ablation and sample final audios for voice conversion from IndicSuperb (Javed et al., 2022), to ensure Indic language coverage and gender-balanced diversity, and from IndicST (Shah et al., 2025), which collates English–Indic open-source audios.
@@ 3.3 Evaluation Framework
iv) Refusal Rate (RR). This is a model safety focused metric which checks if the system declines harmful or unsafe requests instead of executing them. We replicate the implementation in Andriushchenko et al. (2025), using GPT-4o-mini as a semantic judge to classify each response.
@@ 4.1 Models
SpeechLMs. We benchmark 3 SOTA 7B SpeechLMs: (i) KimiAudio 7B (KimiTeam et al., 2025), (ii) Qwen2.5-Omni 7B (Xu et al., 2025), (iii) AudioFlamingo3 7B (Ghosh & Duraiswami, 2025).
ASR-LLMs. In this modular setup, user speech is first transcribed with Whisper v3 Large (Whisperv3), and the text is then passed to an LLM along with tools and instructions. We benchmark three strong LLMs: Qwen-3 8B (Yang et al., 2025a), Gemma3 27B (Team et al., 2025), and LLaMA 3.3 70B (Llama3 70B), enabling comparison between this setup and SpeechLMs.
@@ 5 Limitations and Conclusion
Limitations. We acknowledge following limitations in our work: First, our evaluation does not include speech with background noise, and therefore we do not measure the noise impact on tool call invocation. Second, we do not extend evaluation for multi-turn dialogues for Indic languages, which is also critical for building general-purpose voice assistants. Third, due to prohibitive costs, we exclude closed-source voice assistant systems such as GPT-4o-audio and Gemini-2.5-Pro from our evaluation. Lastly, our study does not evaluate dynamic, real-time tool invocation with interactive user conversation, as explored in frameworks like Yao et al. (2025).
@@
Acknowledgements
We sincerely thank the leadership at Krutrim AI for their support in enabling this research. We also extend our gratitude to the Speech team at Krutrim for their valuable suggestions and insightful discussions that contributed to this work.
@@ Appendix B Additional Related Work: Speech Models
Early audio-language encoders, such as AudioCLIP (Guzhov et al., 2022) and CLAP (Elizalde et al., 2023), learn joint embeddings of speech and text, enabling tasks like cross-modal retrieval, keyword-based speech search, and basic classification. These models primarily focus on representation learning without complex reasoning or generative capabilities. Specialized speech models, including Whisper (Radford et al., 2022), SALM (Chen et al., 2024b), and AudioPALM (Rubenstein et al., 2023), excel in automatic speech recognition (ASR), speech-to-text translation, and speech understanding, enabling transcription, translation, and limited instruction following over speech inputs. Integrated multitask models such as AudioGPT (Huang et al., 2024), WavLLM (Hu et al., 2024), LTU (Gong et al., 2024), and SALMONN (Tang et al., 2024) extend these capabilities to multi-turn dialogue, question answering, and instruction following by combining ASR, speech understanding, and LLM-based reasoning. Recent large audio-language models, including Qwen2-Audio (Chu et al., 2024), KimiAudio 7B (KimiTeam et al., 2025), Qwen2.5-Omni 7B (Xu et al., 2025), and Audio Flamingo 3 (Ghosh & Duraiswami, 2025), further enhance reasoning capabilities over speech, enabling long-form question answering, multi-step instruction execution, and chat-style conversation.
@@ C.1 Density-Based Probabilistic Method
In this method, we start with a set of audio samples from source dataset. Then, each audio sample is passed through an ECAPA-TDNN (Desplanques et al., 2020) model trained on VoxLingua107 to generate fixed-dimensional embeddings that capture both speaker identity and acoustic features:

### v2

@@ 1 Introduction
Advancements in Large Language Models (LLMs) (touvron2023llama; grattafiori2024llama; abdin2025phi; guo2025deepseek; yang2025qwen3) have enabled the development of intelligent agents capable of reasoning (wei2022chain), planning (yao2023react), and executing complex, multi-step tasks through interaction with external tools (qin2024toolllm; patil2024gorilla) and databases (gao2024retrievalaugmentedgenerationlargelanguage). These agentic systems have shown strong performance on tasks such as code generation (rozière2024codellamaopenfoundation; DBLP:journals/corr/abs-2406-11931), document question answering (DBLP:journals/corr/abs-2410-18050), and interactive AI applications, highlighting their potential to automate sophisticated workflows. Most existing research, however, focuses on text-based interactions, overlooking speech as a natural and accessible modality. Extending agents to voice is critical for hands-free, conversational control in real-world applications. Current approaches typically depend on external automatic speech recognition (ASR) models, leaving open the fundamental question of how agents can directly process and respond to spoken input without an intermediate step. SpeechLMs [^1] Although existing models and benchmarks cover multiple formats, including audio, speech, and music, our focus here is specifically on human speech tasks; henceforth we use the term SpeechLMs. capable of instruction following and chat-style interactions (xu2025qwen25omnitechnicalreport; kimiteam2025kimiaudiotechnicalreport) pave the way toward such agents. By avoiding the sequential transcription overhead of ASR–LLM pipelines, SpeechLMs enable lower latency (Appendix F) and natural real-time interactions. They further demonstrate the ability to interpret spoken commands, maintain context across turns, and generate task-relevant responses in natural language, providing a strong foundation for voice assistants.
@@ 2 Related Work
LLM Agent Benchmarks. Interest in evaluating agentic LLMs has grown with advances in their reasoning and decision making capabilities. ToolBench (qin2024toolllm) evaluates models’ ability to invoke external APIs across diverse real-world tasks, while ToolQA (zhuang2023toolqa) assesses LLMs’ use of external tools for question answering via a scalable, automated dataset curation process. Berkeley Function Calling Leaderboard (BFCL) (patil2025the) emphasizes precise API generation across domains and robustness to both single and multiple function calls, and NESTful (basu2025nestfulbenchmarkevaluatingllms) focuses on nested sequences of API calls, where outputs of one call feed into the next. API-Bank and ToolTalk (li2023apibank; farn2023tooltalkevaluatingtoolusageconversational) target multi-turn, dialogue-driven tool-use scenarios, testing sequential API planning and interaction. Tau-bench (yao2025taubench) simulates dynamic conversations with domain-specific tools and policies to evaluate adherence to task rules. AgentHarmBench (andriushchenko2025agentharm) and DoomArena (boisvert2025doomarena) focus on safety and adversarial robustness, testing susceptibility to harmful or unsafe actions. Despite this progress for LLMs, no speech benchmark explicitly evaluates SpeechLMs in such realistic, agentic, and safety-critical settings, underscoring the need for specialized evaluation frameworks.
@@
Speech Datasets and Benchmarks. Large-scale datasets such as LibriSpeech (7178964), CommonVoice (ardila-etal-2020-common), and MuST-C (di-gangi-etal-2019-must) have been foundational for advancing automatic speech recognition (ASR) and speech translation (AST). IndicST (11011192) and Lahaja (javed24_interspeech) extend these tasks to cover diverse Indic speech data. Evaluation suites like SUPERB (yang21c_interspeech) and SLUE (shon-etal-2023-slue) standardize the assessment of tasks such as intent classification, named entity recognition, and slot filling, with IndicSUPERB (javed2022indicsuperbspeechprocessinguniversal) further supporting Indic languages. However, these benchmarks primarily target simpler tasks like transcription, translation, NER and do not assess reasoning or decision-making over spoken content. To address this gap, recent work has begun exploring reasoning in the audio domain. Audio-CoT (ma2025audiocotexploringchainofthoughtreasoning) introduces chain-of-thought prompting for structured multistep inference on speech input, while MMAU (sakshi2025mmau) provides a large-scale benchmark of 10k audio clips covering 27 reasoning skills, such as temporal reasoning and causal inference, in speech, music, and environmental sounds. AIR-Bench (yang-etal-2024-air) and AudioBench (DBLP:journals/corr/abs-2406-16020) extend the scope to open-ended instruction following on various types of audio and speech, whereas VoiceBench (DBLP:journals/corr/abs-2410-17196) emphasizes robustness and generalization by converting text instruction into spoken form with real-world noise and speaker variation. More recently, SpeechR (yang2025speechrbenchmarkspeechreasoning) directly targets high-level reasoning on speech, focusing on logical deduction, and commonsense problem solving. We also provide an extended discussion of related work on speech models in Appendix B.
@@ 3.2.1 Tool Sourcing
Single Tool, Single Tool with Retrieval, and Parallel Tool Invocation. For these categories, we leverage BFCL (patil2025the), which provides well-structured functions for such cases. Single Tool Call directly adapts the simple tool subset from BFCL. The Single Tool with Retrieval category uses the multiple tool subset, requiring the model to select the most relevant tool and fill the arguments. The Parallel Tool category leverages the parallel multiple subset, where multiple independent tool calls are invoked simultaneously. In total, we incorporate 458 functions here.
Sequential Dependent Tools. While benchmarks like NESTful (basu2025nestfulbenchmarkevaluatingllms) focus on sequential tool invocation in specialized domains such as mathematics and coding, they do not capture the practical, everyday tasks expected of real-world voice assistants. These tasks often require chaining interdependent tools to complete workflows such as booking a cab, ordering food, or managing payments. To address this, we designed a set of 21 tools across three realistic agents: i) Cab booking, ii) Food ordering, and iii) Payment services. The toolsets are presented in Appendix H.
Multi-Turn Dialogue Tools. For dialogue-based tool invocation, we adopt tools from the Level-2 subset of API-Bank (li2023apibank), which are designed to support Retrieval+Call based on multi-turn user interactions. Using this subset, we incorporate a total of 49 tools for this category, enabling evaluation of conversationally grounded, multi-turn tool-calling capabilities.
Safety Evaluation Tools. For evaluating the safety of agentic behavior, we utilize tools from the AgentHarm (andriushchenko2025agentharm). It encompasses 11 harm categories, including fraud, cybercrime, and harassment. These tasks are designed to assess whether models can refuse harmful agentic requests. By integrating these tools, we enable comprehensive evaluation of an agent’s robustness against adversarial and unsafe tool usage scenarios.
@@ 3.2.2 Text Query & Dialogue Generation
ii). Sequential Dependent Tool Calling (SeqDep): To capture realistic use cases, we design three service agents: cab booking, food ordering, and payments by creating seed data with available tools and expected outputs. Using GPT-4o-mini (hurst2024gpt), we generate queries conditioned on tool schemas and dependency chains, ensuring multiple interdependent calls (e.g., booking a cab after retrieving location coordinates or completing a food order with stored address and payment details). This yields 40 complex queries each in English and Hindi, paired with tool-call responses, capturing dependency-driven tasks beyond simple or parallel invocation.
@@ 3.2.3 Model Instructions and Indic Multilingual Extension
To standardize behavior across models, we design category-specific system instructions that direct the model to produce tool calls strictly in Python syntax, following patil2025the. This prevents free-form or mixed response which cannot be deterministically parsed. To further anchor the model’s behavior, we provide a one-shot example in the instruction that clarifies the output format without imposing strong task-specific biases. For multilingual settings, we additionally instruct the model to generate tool calls exclusively in English, aligning with real-world usage scenarios. Additionally, for safety evaluations we append a refusal prompt in the instructions (examples in Appendix J.1).
We extend VAB to five Indic languages by translating the English Indian context queries into the target languages. For Malayalam, we employ Llama-3.3 70B, while for the remaining languages we use Gemma3 27B, this is based on human eval results shown in (2025bhasha). We complement this with a lightweight quality control pipeline that flags issues such as script mixing and unknown tokens. The validated queries are then fed into the diversity based TTS pipeline.
@@ 3.2.4 Diversity Based TTS Generation
In synthetic speech generation, the absence of real speakers and natural voices necessitates methods that ensure diversity in the generated data, motivating new selection strategies to build robust and representative benchmarks. Following IndicSynth (sharma-etal-2025-indicsynth), which employed the VoxLingua107 ECAPA-TDNN model (Desplanques_2020), trained on diverse Indic languages and accents, to evaluate the linguistic authenticity of synthetic speech, we adopt ECAPA-TDNN embeddings for our diversity analysis. Adapting diversity principles from word embedding literature, we transfer these ideas to the audio domain using ECAPA-TDNN embeddings. Specifically, we ablate three strategies for selecting maximally diverse audio samples: i) Determinantal Point Processes (DPP) (wang2025diversity), ii) Farthest Point Sampling (FPS) adapted from PointNet++ (PointNet++), and iii) a Density-based Probabilistic Method (Appendix C.1). Diversity is quantified using the mean distance to the nearest selected point (yang2025measuringdatadiversityinstruction), a metric that captures coverage of the embedding space. Our evaluation (Appendix C.2) shows that FPS (Algorithm 1) consistently achieves the highest diversity scores on our dataset, establishing it as the most effective strategy. We conduct this ablation and sample final audios for voice conversion from IndicSuperb (javed2022indicsuperbspeechprocessinguniversal), to ensure Indic language coverage and gender-balanced diversity, and from IndicST (11011192), which collates English–Indic open-source audios.
@@ 3.3 Evaluation Framework
iv) Refusal Rate (RR). This is a model safety focused metric which checks if the system declines harmful or unsafe requests instead of executing them. We replicate the implementation in andriushchenko2025agentharm, using GPT-4o-mini as a semantic judge to classify each response.
@@ 4.1 Models
SpeechLMs. We benchmark 3 SOTA 7B SpeechLMs: (i) KimiAudio 7B (kimiteam2025kimiaudiotechnicalreport), (ii) Qwen2.5-Omni 7B (xu2025qwen25omnitechnicalreport), (iii) AudioFlamingo3 7B (ghosh2025audio).
ASR-LLMs. In this modular setup, user speech is first transcribed with Whisper v3 Large (Whisperv3), and the text is then passed to an LLM along with tools and instructions. We benchmark three strong LLMs: Qwen-3 8B (yang2025qwen3technicalreport), Gemma3 27B (gemmateam2025gemma3technicalreport), and LLaMA 3.3 70B (Llama3 70B), enabling comparison between this setup and SpeechLMs.
@@ 5 Limitations and Conclusion
Limitations. We acknowledge following limitations in our work: First, our evaluation does not include speech with background noise, and therefore we do not measure the noise impact on tool call invocation. Second, we do not extend evaluation for multi-turn dialogues for Indic languages, which is also critical for building general-purpose voice assistants. Third, due to prohibitive costs, we exclude closed-source voice assistant systems such as GPT-4o-audio and Gemini-2.5-Pro from our evaluation. Lastly, our study does not evaluate dynamic, real-time tool invocation with interactive user conversation, as explored in frameworks like yao2025taubench.
@@ Appendix B Additional Related Work: Speech Models
Early audio-language encoders, such as AudioCLIP (9747631) and CLAP (10095889), learn joint embeddings of speech and text, enabling tasks like cross-modal retrieval, keyword-based speech search, and basic classification. These models primarily focus on representation learning without complex reasoning or generative capabilities. Specialized speech models, including Whisper (radford2022robustspeechrecognitionlargescale), SALM (chen2024salm), and AudioPALM (rubenstein2023audiopalmlargelanguagemodel), excel in automatic speech recognition (ASR), speech-to-text translation, and speech understanding, enabling transcription, translation, and limited instruction following over speech inputs. Integrated multitask models such as AudioGPT (DBLP:conf/aaai/HuangLYSCYWHHLR24), WavLLM (DBLP:journals/corr/abs-2404-00656), LTU (gong2024listen), and SALMONN (tang2024salmonn) extend these capabilities to multi-turn dialogue, question answering, and instruction following by combining ASR, speech understanding, and LLM-based reasoning. Recent large audio-language models, including Qwen2-Audio (chu2024qwen2audiotechnicalreport), KimiAudio 7B (kimiteam2025kimiaudiotechnicalreport), Qwen2.5-Omni 7B (xu2025qwen25omnitechnicalreport), and Audio Flamingo 3 (ghosh2025audio), further enhance reasoning capabilities over speech, enabling long-form question answering, multi-step instruction execution, and chat-style conversation.
@@ C.1 Density-Based Probabilistic Method
In this method, we start with a set of audio samples from source dataset. Then, each audio sample is passed through an ECAPA-TDNN (Desplanques_2020) model trained on VoxLingua107 to generate fixed-dimensional embeddings that capture both speaker identity and acoustic features: